# Subexponential lower bounds for $f$-ergodic Markov processes

### v1

@@ Abstract.
We provide a criterion for establishing lower bounds on the rate of convergence in $f$-variations of an ergodic Markov processes to its invariant measure. The criterion consists of super and submartingale conditions for certain functionals of the Markov process. Our results complement the existing literature on the stability of Markov processes, based on Lyapunov drift conditions, which primarily focuses on upper bounds. We apply our theory to elliptic diffusions, Lévy-driven Ornstein-Uhlenbeck processes and hypoelliptic stochastic Hamiltonian models with known polynomial/stretched exponential upper bounds on their rates of convergence. Our lower bounds match asymptotically the known upper bounds for these classes of models, establishing their rate of convergence to stationarity.
Key words: subgeometric ergodicity; lower bounds; rate of convergence in f-variation and total variation; invariant measure; return times; Lyapunov functions; Foster type drift conditions.
@@ 1. Introduction
To the best of our knowledge, these are the first result to provide general Foster-Lyapunov-type drift conditions for lower bounds on the convergence rates of ergodic Markov processes. The super and submartingale properties of the processes in (1) are typically verified using the infinitesimal characteristics of $X$ and Itô’s formula. In conjunction with the classical drift conditions for upper bounds [9, 12, 24], our results offer a comprehensive and robust approach to quantifying the convergence rate of ergodic Markov processes.
As pointed out in [13, Sec. 5], deriving a lower bound on the rate of convergence to the invariant distribution typically requires a deep understanding of the tail of the invariant measure, which is available only in very special cases (e.g. Langevin diffusions). The main contribution of this paper consists of providing a robust framework, which enables us to derive lower bounds on the convergence rates without prior knowledge of the tail behavior of the invariant measure. In the context of stochastic differential equations, our theory makes it possible to establish lower bounds for general equations with multiplicative noise, in contrast to being limited to special cases such as Langevin-type equations with additive noise [13, Sec. 7.1] (see [3] for short YouTube presentations describing the applications and proofs of our main results).
The remainder of the paper is structured as follows. Section 2 states the novel $\mathbf{L}$-drift condition and formulates our main theorems for ergodic Markov processes. Section 3 presents the applications of the results of Section 2 to classes of models, studied in [9, 12], exhibiting subexponential (and exponential) ergodicity. Sections 4 and 5 contain the proofs of the results of Section 2. More precisely, Sections 4 states and proves lower bounds for the tails of return times to bounded sets for càdlàg semimartingales. Section 5 applies these results in the context of càdlàg ergodic Markov processes to prove the theorems of Section 2. The proofs of the results, stated in Section 3, for ergodic elliptic diffusions, Lévy-driven Ornstein-Uhlenbeck processes and hypoelliptic stochastic damping Hamiltonian system are in Section 6. Finally, Section 7 offers concluding remarks, highlights some open questions and describes future directions of research.
@@ 2.2. Lower bounds: the tails of $\pi$ and $f$-variation
(ii) Let $\Psi:[1,\infty)\to[1,\infty)$ be a differentiable, increasing, submultiplicative [^2] A function $\Psi:[1,\infty)\to[1,\infty)$ is submultiplicative if it satisfies $\Psi(r_{1}+r_{2})\leq C\Psi(r_{1})\Psi(r_{2})$ for some constant $C\in(0,\infty)$ and all $r_{1},r_{2}\in[1,\infty)$ [31, Def. 25.2]. function satisfying the following: for any $\ell\in(\ell_{0},\infty)$, there exists a constant $C_{\ell}\in(0,\infty)$ such that
@@
Let Assumption  (L($V$,$\varphi$,$\Psi$)). hold and set $f_{\star}\equiv 1$. Assume there exists continuous functions $h,a:[1,\infty)\to{\mathbb{R}}_{+}$ and $v:{\mathcal{X}}\times{\mathbb{R}}_{+}\to[1,\infty)$, satisfying conditions (a) and (b) in Theorem 2.2. Let $r_{1}:{\mathcal{X}}\times[1,\infty)\to{\mathbb{R}}_{+}$, $r_{1}\coloneqq a\circ A^{-1}\circ(2v)$, be as in Theorem 2.2. Then the following lower bound holds
@@
    The submultiplicative function [31, Def. 25.2] (cf. footnote on page 2) $\Psi$ in  (L($V$,$\varphi$,$\Psi$)). transforms the process $V(X)$ into a submartingale. The function $\Psi$ features in lower bounds in our theorems through definition (3). Since a product of submultiplicative functions is submultiplicative and, by [31, Prop. 25.4], $r\mapsto g(cr+\gamma)^{\alpha}$ (where $c,\gamma,\alpha>0$) is submultiplicative if $g$ is, it follows that $r\mapsto r^{a}(\log r)^{d}\exp(br^{p})$ is submultiplicative for (i) $b>0$ and $p\in(0,1)$, (ii) $b=0$ and $a>0$ or (iii) $b=a=0$ and $d\geq 0$. This form is similar to the subgeometric rate functions used in [9, 12]. In fact, the rate functions in [9, 12] are submultiplicative [32, Eq. (5)], a crucial fact used in the proofs of [9, 12].
@@ 2.3. Modulated moments for the process
For a set $D\in{\mathcal{B}}({\mathcal{X}})$ and $\delta\in(0,\infty)$, let $\tau_{D}(\delta)\coloneqq\inf\{t>\delta:X_{t}\in D\}$ (with convention $\inf\emptyset\coloneqq\infty$) be the first hitting time of $D$ after time $\delta$ (recall that $\tau_{D}(\delta)$ is an $({\mathcal{F}}_{t})$-stopping time by [15, Thm 1.27]). In Theorem 2.5 we provide, under the L-drift condition  (L($V$,$\varphi$,$\Psi$))., lower bounds on the tails of return times to arbitrary subsets of sublevel sets of the Lyapunov function $V$. Crucially, in Section 5.2 below, we prove that, under the L-drift condition, all petite sets are contained in the sublevel sets of $V$, making Theorem 2.5 applicable to any petite set of $X$.
@@ 2.4.1. Generators and drift conditions
The classical Lyapunov drift condition [9, 12, 24] requires the process $V(X)$ to satisfy the supermartingale condition in (10) below. If such $V$ is available, verifying  (L($V$,$\varphi$,$\Psi$)). reduces to finding scalar functions $\varphi$ and $\Psi$, such that $1/V(X)-\int_{0}^{\cdot}\varphi\circ(1/V)(X_{s}){\mathrm{d}}s$ and $\Psi\circ V(X)$ are a super and a submartingale, respectively. Since $\Psi$ is increasing, the Lyapunov function $V:{\mathcal{X}}\to[1,\infty)$ determines the level sets of $\Psi\circ V$, while $\Psi:[1,\infty)\to[1,\infty)$ modulates only the growth of $\Psi\circ V$. Thus, identifying $\Psi$ is typically straightforward if $V$ is given and the task is to find the slowest growing $\Psi$ so that $\Psi\circ V(X)$ is a submartingale. Identifying $\varphi$ for a given $V$ is also typically a simple task, as it reduces to bounding the drift of $1/V(X)$ using Theorem 2.7(a) in Subsection 2.4.1.
@@
Implication (II) in Table 1 above uses the fact that, under  (L($V$,$\varphi$,$\Psi$))., all petite sets of $X$ are bounded and, crucially, that the lower bounds on the probability $\mathbb{P}_{x}\left(\int_{0}^{\tau_{D}(\delta)}h\circ V(X_{s}){\mathrm{d}}s\geq r\right)$ hold for all non-decreasing functions $h:[1,\infty)\to[1,\infty)$. These facts, together with the well-known characterisation of Meyn and Tweedie in [25, Thm 1.2(b)] of the integrability with respect to $\pi$, yield a lower bound on the tail of the invariant measure $\pi$ in Theorem 2.1.
Implication (III) in Table 1 above requires a comparison of the lower bound on the tail of the invariant measure $\pi$, obtained via implication (II), with the upper bound on the tail of the law of $V(X_{t})$, controlled by the expected growth of the process $V(X)$. Finally, an application of Lemma 5.1, which is generalisation to $f$-variation norms of a result in Hairer’s notes [13, Thm 5.1], yields our lower bound on the rate of convergence (Theorem 2.2).
@@ 2.6. Related literature
Since this approach relies on transforming the state space with a Lyapunov function $V$, key information may be lost, potentially resulting in poor upper bound estimates on the convergence rate (see the motivating example in Section 3.1 below; see also Example 3.5). This naturally motivates a general study of lower bounds in the context of Lyapunov functions presented in this paper, particularly in the case when the upper bounds on the rate of convergence are subgeometric. To the best of our knowledge, in the context of Lyapunov functions, the present paper is the first work dedicated to this problem. As explained in Section 2.4, our results naturally augment the existing Lyapunov function approach for the stability of Markov processes and provide a robust method for checking the quality of upper bound estimates obtained via a given Lyapunov function.
@@
A potential theoretic approach for the stability of a class of Lévy-driven OU-processes, arising as limits of multiclass many-server queues, has been developed in [1, 29]. In contrast to our results, the lower bounds [29, Thm 1.2] for the general case of their models hold along a sequence of times $t_{n}\to\infty$. The theory in [29] gives no information on the sparsity of the sequence $(t_{n})_{n\in{\mathbb{N}}}$. However, in some special cases of the models in [1, 29], matching lower and upper bounds for all $t\in[1,\infty)$ are established using Hairer’s result [13, Thm 5.1]. In these special cases, the precise decay of the tails of the invariant measure of the model in [1] is established using analytical methods.
@@
We now apply the results of Section 2 to stochastic models used across probability and its application. We mostly consider models where Lyapunov drift conditions have been developed to obtain upper bounds on the rates of convergence. The examples are chosen to demonstrate the robustness and the complementary nature of our approach to the existing theory for upper bounds and cover the classes of models studied in [9, Sec. 5] and [12, Sec. 3]. In particular, we analyse models exhibiting polynomial, stretched exponential and exponential ergodicity and in all cases provide lower bounds, which asymptotically match known upper bounds. More precisely, we give lower bounds in $f$-variation for polynomially ergodic elliptic and hypoelliptic diffusions, studied in [12, Sec. 3] and [9, 34], respectively. For brevity, in all other examples we consider total variation (even though our methods could handle $f$-variation in these cases) and show that our lower bounds match asymptotically the upper bounds in [12, 9].

### v2

@@ Abstract.
We provide a criterion for establishing lower bounds on the rate of convergence in $f$-variation of a continuous-time ergodic Markov process to its invariant measure. The criterion consists of novel super- and submartingale conditions for certain functionals of the Markov process. It provides a general approach for proving lower bounds on the tails of the invariant measure and the rate of convergence in $f$-variation of a Markov process, analogous to the widely used Lyapunov drift conditions for upper bounds. Our key technical innovation produces lower bounds on the tails of the heights and durations of the excursions from bounded sets of a continuous-time Markov process using path-wise arguments.
We apply our theory to elliptic diffusions and Lévy-driven stochastic differential equations with known polynomial/stretched exponential upper bounds on their rates of convergence. Our lower bounds match asymptotically the known upper bounds for these classes of models, thus establishing their rate of convergence to stationarity. The generality of the approach suggests that, analogous to the Lyapunov drift conditions for upper bounds, our methods can be expected to find applications in many other settings.
Key words: subgeometric ergodicity; lower bounds; rate of convergence in f-variation and total variation; invariant measure; return times; Lyapunov functions and Foster-type drift conditions.
@@ 1. Introduction
Building on ideas in [15, 13], we provide general Foster-Lyapunov-type drift conditions implying lower bounds for all times on the convergence rates of ergodic Markov processes. The super- and submartingale properties of the processes in (1) are typically verified using the infinitesimal characteristics of $X$ and Itô’s formula. In conjunction with the classical drift conditions for upper bounds [9, 12, 27], our results offer a comprehensive and robust approach to quantifying the convergence rate of ergodic Markov processes.
As pointed out in [13, Sec. 5], deriving a lower bound on the rate of convergence to stationarity typically requires a deep understanding of the tail of the invariant measure, which is available only in very special cases (e.g. Langevin diffusions). The main contribution of this paper consists of providing a robust framework for deriving lower bounds on the convergence rates without prior knowledge of the tail behavior of the invariant measure. For example, in the context of (possibly Lévy-driven) stochastic differential equations with multiplicative noise, our theory makes it possible to establish lower bounds on $f$-variation for all times (as opposed to along a sparse sequence of times going to infinity as in [15]). To the best of our knowledge, such results were previously only available for Langevin diffusions [15, Thm 2.1], cf. Section 2.6 below.
The remainder of the paper is structured as follows. Section 2 states the novel $\mathbf{L}$-drift condition and formulates our main theorems for ergodic Markov processes. Section 3 presents the applications of the results of Section 2 to classes of models, studied in [9, 12], exhibiting subexponential (and exponential) ergodicity. Sections 4 and 5 contain the proofs of the results of Section 2. More precisely, Sections 4 states and proves lower bounds for the tails of return times to bounded sets for càdlàg semimartingales. Section 5 applies these results in the context of càdlàg ergodic Markov processes to prove the theorems of Section 2. The proofs of the results, stated in Section 3, for ergodic elliptic diffusions, Lévy-driven stochastic differential equations and hypoelliptic stochastic damping Hamiltonian system are in Section 6. Finally, Section 7 offers concluding remarks, highlights some open questions and describes future directions of research. (See [3] for short YouTube presentations describing the applications and proofs of our main results.)
@@ 2.2. Lower bounds: the tails of $\pi$ and $f$-variation
(ii) Let $\Psi:[1,\infty)\to[1,\infty)$ be a differentiable, increasing, submultiplicative [^2] A function $\Psi:[1,\infty)\to[1,\infty)$ is submultiplicative if it satisfies $\Psi(r_{1}+r_{2})\leq C\Psi(r_{1})\Psi(r_{2})$ for some constant $C\in(0,\infty)$ and all $r_{1},r_{2}\in[1,\infty)$ [33, Def. 25.2]. function satisfying the following: for any $\ell\in(\ell_{0},\infty)$, there exists a constant $C_{\ell}\in(0,\infty)$ such that
@@
Let Assumption  (L($V$,$\varphi$,$\Psi$)). hold and set $f_{\star}\equiv 1$. Assume there exist continuous functions $h,a:[1,\infty)\to{\mathbb{R}}_{+}$ and $v:{\mathcal{X}}\times{\mathbb{R}}_{+}\to[1,\infty)$, satisfying conditions (a) and (b) in Theorem 2.2. Let $r_{1}:{\mathcal{X}}\times[1,\infty)\to{\mathbb{R}}_{+}$, $r_{1}\coloneqq a\circ A^{-1}\circ(2v)$, be as in Theorem 2.2. Then the following lower bound holds
@@
    The submultiplicative function [33, Def. 25.2] (cf. footnote on page 2) $\Psi$ in  (L($V$,$\varphi$,$\Psi$)). transforms the process $V(X)$ into a submartingale. The function $\Psi$ features in lower bounds in our theorems through definition (3). Since a product of submultiplicative functions is submultiplicative and, by [33, Prop. 25.4], $r\mapsto g(cr+\gamma)^{\alpha}$ (where $c,\gamma,\alpha>0$) is submultiplicative if $g$ is, it follows that $r\mapsto r^{a}(\log r)^{d}\exp(br^{p})$ is submultiplicative for (i) $b>0$ and $p\in(0,1)$, (ii) $b=0$ and $a>0$ or (iii) $b=a=0$ and $d\geq 0$. This form is similar to the subgeometric rate functions used in [9, 12]. In fact, the rate functions in [9, 12] are submultiplicative [34, Eq. (5)], a crucial fact used in the proofs of [9, 12].
@@ 2.3. Modulated moments for the process
For a set $D\in{\mathcal{B}}({\mathcal{X}})$ and $\delta\in(0,\infty)$, let $\tau_{D}(\delta)\coloneqq\inf\{t>\delta:X_{t}\in D\}$ (with convention $\inf\emptyset\coloneqq\infty$) be the first hitting time of $D$ after time $\delta$ (recall that $\tau_{D}(\delta)$ is an $({\mathcal{F}}_{t})$-stopping time by [16, Thm 1.27]). In Theorem 2.5 we provide, under the L-drift condition  (L($V$,$\varphi$,$\Psi$))., lower bounds on the tails of return times to arbitrary subsets of sublevel sets of the Lyapunov function $V$. Crucially, in Section 5.2 below, we prove that, under the L-drift condition, all petite sets are contained in the sublevel sets of $V$, making Theorem 2.5 applicable to any petite set of $X$.
@@ 2.4.1. Generators and drift conditions
The classical Lyapunov drift condition [9, 12, 27] requires the process $V(X)$ to satisfy the supermartingale condition in (10) below. If such $V$ is available, verifying  (L($V$,$\varphi$,$\Psi$)). reduces to finding scalar functions $\varphi$ and $\Psi$, such that $1/V(X)-\int_{0}^{\cdot}\varphi\circ(1/V)(X_{s}){\mathrm{d}}s$ and $\Psi\circ V(X)$ are a super- and a submartingale, respectively. Since $\Psi$ is increasing, the Lyapunov function $V:{\mathcal{X}}\to[1,\infty)$ determines the level sets of $\Psi\circ V$, while $\Psi:[1,\infty)\to[1,\infty)$ modulates only the growth of $\Psi\circ V$. Thus, identifying $\Psi$ is typically straightforward if $V$ is given and the task is to find the slowest growing $\Psi$ so that $\Psi\circ V(X)$ is a submartingale. Identifying $\varphi$ for a given $V$ is also typically a simple task, as it reduces to bounding the drift of $1/V(X)$ using Theorem 2.7(a) in Subsection 2.4.1.
@@
Implication (II) in Table 1 above uses the fact that, under  (L($V$,$\varphi$,$\Psi$))., all petite sets of $X$ are bounded and, crucially, that the lower bounds on the probability $\mathbb{P}_{x}\left(\int_{0}^{\tau_{D}(\delta)}h\circ V(X_{s}){\mathrm{d}}s\geq r\right)$ hold for all non-decreasing functions $h:[1,\infty)\to[1,\infty)$. These facts, together with the well-known characterisation of Meyn and Tweedie in [24, Thm 1.2(b)] of the integrability with respect to $\pi$, yield a lower bound on the tail of the invariant measure $\pi$ in Theorem 2.1.
Implication (III) in Table 1 above requires a comparison of the lower bound on the tail of the invariant measure $\pi$, obtained via implication (II), with the upper bound on the tail of the law of $V(X_{t})$, controlled by the expected growth of the process $V(X)$. Finally, an application of Lemma 5.1, which is generalisation to $f$-variation norms of [15, Thm 3.6] (see also [13, Thm 5.1]), yields our lower bound on the rate of convergence in Theorem 2.2.
@@ 2.6. Related literature
Lyapunov functions and lower bounds.
A general Lyapunov-function approach to lower bounds on the total variation distance along a sequence of times tending to infinity is developed in [15]. A key step [15, Thm 3.6] consists of converting lower bounds on the tails of the invariant measure to the lower bounds on the total variation, an idea also exploited in the present paper (see its generalisation from total to $f$-variation in Lemma 5.1 below). Crucially, in [15, Thm 3.2], Hairer gives criteria (based on ideas from [37]) for establishing lower bounds on total variation along a sequence of times tending to infinity when the invariant measure is not explicitly known. These criteria can be summarised briefly as follows: let $W_{1}$ and $W_{2}$ be $C^{2}$-functions on ${\mathbb{R}}^{n}$ such that $W_{1}/W_{2}\to 0$ as $W_{1}\to\infty$ and the inequalities ${\mathcal{A}}W_{1}\geq 0$ and ${\mathcal{A}}W_{2}\leq F$ hold outside a compact set for some function $F$, where ${\mathcal{A}}$ denotes the generator of $X$. Under these conditions [15, Thm 3.2] implies $\int F{\mathrm{d}}\pi=\infty$ and, together with [15, Thm 3.6], yields a sequence of times $t_{n}\to\infty$ at which a lower bound on the total variation (based on non-integrability of $F$) holds.
The critical step in [15] consists of finding a non-integrable function $F$, which essentially amounts to the function ${\mathcal{A}}W_{2}$ being non-integrable: $\int{\mathcal{A}}W_{2}{\mathrm{d}}\pi=\infty$. Consequently, the lower bounds on the tails of $\pi$ hold only along a (possibly sparse) sequence of levels, yielding a (possibly sparse) sequence of times $t_{n}\to\infty$ at which the lower bound on the total variation can be established. In contrast, our approach provides lower bounds for all times by essentially analysing lower bounds as upper bounds of $1/V(X)$ (as usual $V$ denotes a Lypuanov function for $X$). In the language of [15] this involves estimating the drift ${\mathcal{A}}(1/V)$, and then using path-wise arguments for semimartingales (Lemma 4.2 below) to characterise $\int h\circ V{\mathrm{d}}\pi=\infty$ for all increasing scalar functions $h$. This idea allows us to establish lower bounds on the tails of $\pi$ for all levels and hence get lower bounds on the convergence rate in $f$-variation for all times. Moreover, this suggests that establishing lower bounds for all times using the approach in [15] directly would at the very least require a sufficiently rich family of functions $W_{2}^{(\kappa)}$, satisfying the assumptions of [15, Thm 3.2] (and hence $\int{\mathcal{A}}W_{2}^{(\kappa)}{\mathrm{d}}\pi=\infty$) for all values of the parameter $\kappa$.
In the context of elliptic diffusions studied in Section 3.1 below, our results may be viewed as a generalisation of the bounds along $t_{n}\to\infty$, obtained via [15, Thm 3.2], to all times $t\in[1,\infty)$. If the process $X$ has jumps with heavy tails as in Section 3.2 below, [15, Thm 3.2] appears to be difficult to apply: for the Lévy-driven stochastic differential equation of Section 3.2, $W_{1}$ is either not integrable with respect to the jump measure of $X$ or, if it is, we have ${\mathcal{A}}W_{1}<0$ outside of a compact (see the formula for ${\mathcal{A}}$ in equation (42) and Remark 6.6 below for details). This makes the condition $0\leq{\mathcal{A}}W_{1}<\infty$ hard to satisfy outside of any compact set. Finally, it is feasible that our methods could yield novel insights for the hypoelliptic Hamiltonian system studied in [15] and possibly extend the result on lower bounds in [15, Thm 1.1] to all times. The function $W_{1}$ in [15, eq. (5.15)], satisfying ${\mathcal{A}}W_{1}\geq 0$ and thus making $W_{1}(X)$ a submartingale, is a natural candidate for $\Psi\circ V$ in the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$)).Assumption. Determining the function $V$ and estimating the asymptotic behaviour of ${\mathcal{A}}(1/V)$, so that  (L($V$,$\varphi$,$\Psi$)).Assumption holds, would enable the application of our results. This is left for future research.
@@
Since this approach relies on transforming the state space with a Lyapunov function $V$, key information may be lost, potentially resulting in poor upper bound estimates on the convergence rate (see the motivating example in Section 3.1 below; see also Example 3.5). This naturally motivates a general study of lower bounds in the context of Lyapunov functions presented in this paper, particularly in the case when the upper bounds on the rate of convergence are subexponential. As explained in Section 2.4, our results naturally augment the existing Lyapunov function approach for the stability of Markov processes and provide a robust method for checking the quality of upper bound estimates obtained via a given Lyapunov function.
@@
A potential theoretic approach for the stability of a class of Lévy-driven OU-processes, arising as limits of multiclass many-server queues, has been developed in [1, 31]. In contrast to our results, the lower bounds [31, Thm 1.2] for the general case of their models hold along a sequence of times $t_{n}\to\infty$. The theory in [31] gives no information on the sparsity of the sequence $(t_{n})_{n\in{\mathbb{N}}}$. However, in some special cases of the models in [1, 31], matching lower and upper bounds for all $t\in[1,\infty)$ are established using Hairer’s result [13, Thm 5.1]. In these special cases, the precise decay of the tails of the invariant measure of the model in [1] is established using analytical methods.
Mixing coefficients of Rosenblatt and Kolmogorov
Convergence to stationarity of an ergodic Markov process can also be quantified via Rosenblatt’s and Kolmogorov’s mixing coefficients. Veretennikov [35] provides upper bounds on the coefficients for polynomialy ergodic elliptic diffusions, cf. Section 3.1.1 below. In the special case when the noise is additive, the corresponding lower bounds on the coefficients are given in [36] (for discrete-time additive-noise case see [18]).
@@
We now apply the results of Section 2 to stochastic models used across probability and its application. We mostly consider models where Lyapunov drift conditions have been developed to obtain upper bounds on the rates of convergence. The examples are chosen to demonstrate the robustness and the complementary nature of our approach to the existing theory for upper bounds: we cover the classes of models studied in [9, Sec. 5] and [12, Sec. 3]. In particular, we analyse models exhibiting polynomial, stretched exponential and exponential ergodicity and in all cases provide lower bounds, which asymptotically match known upper bounds. More precisely, we give lower bounds in $f$-variation for polynomially ergodic elliptic and hypoelliptic diffusions, studied in [12, Sec. 3] and [9, 38], respectively. For brevity, in all other examples we consider total variation (even though our methods could handle $f$-variation in these cases) and show that our lower bounds match asymptotically the upper bounds in [12, 9].